Just before a referendum on a school​ budget, a local newspaper polls

436

voters to predict whether the budget will pass. Suppose the budget has the support of

55%

of the voters. What is the probability that the​ newspaper's sample will lead it to predict​ defeat?
Answer #1

p=0.55
q=1-0.55=0.45
n=436

P(Sample will lead to predict defeat)

= P(p < 0.50)

= P(z < (0.50 - 0.55)/sqrt(0.55*0.45/436))

= P(z < (0.50 - 0.55)/0.0238)

= P(z < -2.0985)

= 0.0179305

.........................

Kindly give thumbs up if you like answer.